OpenStudy (anonymous):

why does tan 60 degrees equal the square root of 3? If I use a calculator to calculate I get 1.7 but how would I guess square root of 3?

OpenStudy (anonymous):

\[\tan 60=\frac{\sin 60}{\cos 60}=\frac{\frac{\sqrt{3}}{2}}{\frac{1}{2}}\] You sure your calculator's using degrees instead of radians?

OpenStudy (anonymous):

Okay, according to WolframAlpha: http://www.wolframalpha.com/input/?i=tan+%2860%29 \[\sqrt{3} \approx 1.732...\]

OpenStudy (anonymous):

oh there is a whole formula? The book just introduced them... m = tan = dy/dx... ok must look into tan/cos/sin then..

OpenStudy (anonymous):

got the answer but the back says square root of 3 instead of 1.7 so okay thanks it is not a simple answer...

OpenStudy (asnaseer):

you can derive the answer by looking at an equilateral triangle:  in this triangle, the height 'h' of the triangle is found using Pythagorus as:\[h=\sqrt{x^2-\frac{x^2}{4}}=\sqrt{\frac{3x^2}{4}}=\frac{x\sqrt{3}}{2}\]so we can see then that:\[\tan(60)=\frac{h}{\frac{x}{2}}=\frac{\frac{x\sqrt{3}}{2}}{\frac{x}{2}}=\sqrt{3}\]